# Fine-tuning can Help Detect Pretraining Data from Large Language Models

### v1

@@ Abstract
In the era of large language models (LLMs), detecting pretraining data has been increasingly important due to concerns about fair evaluation and ethical risks. Current methods differentiate members and non-members by designing scoring functions, like Perplexity and Min-k%. However, the diversity and complexity of training data magnifies the difficulty of distinguishing, leading to suboptimal performance in detecting pretraining data. In this paper, we first explore the benefits of unseen data, which can be easily collected after the release of the LLM. We find that the perplexities of LLMs perform differently for members and non-members, after fine-tuning with a small amount of previously unseen data. In light of this, we introduce a novel and effective method termed Fine-tuned Score Deviation (FSD), which improves the performance of current scoring functions for pretraining data detection. In particular, we propose to measure the deviation distance of current scores after fine-tuning on a small amount of unseen data within the same domain. In effect, using a few unseen data can largely decrease the scores of all non-members, leading to a larger deviation distance than members. Extensive experiments demonstrate the effectiveness of our method, significantly improving the AUC score on common benchmark datasets across various models.
@@ 1 Introduction
In the literature, current works of detecting pretraining data primarily focus on designing scoring functions to differentiate members (i.e., seen data during pretraining) and non-members (unseen). For example, previous work shows that sequences leak in the training data tend to have lower perplexity (i.e., higher likelihood) than non-members (Li, 2023). Min-kleverages the kof tokens with minimum token probabilities of a text for detection, assuming that trained data tends to contain fewer outlier tokens (Shi et al., 2024). However, non-member data can obtain low perplexities by including frequent or repetitive texts, while members may contain rare tokens that result in high perplexities. This casts significant doubt on utilizing those scoring functions for detecting pretraining data. Consequently, this issue prompts us to present a preliminary attempt to enlarge the difference between members and non-members for pretraining datasets of LLMs.
@@
Therefore, our key idea behind FSD is to enlarge the gap between members and non-members by exposing the LLM to a few non-members. This can be accomplished by measuring the deviation distance of current scores (See Figure 1), owing to the self-supervised fine-tuning on a few non-members. In effect, the fine-tuning largely decreases the scores of non-member data, resulting in more distinguishable seen and unseen data. In practical applications, it is easy to collect a small amount of unseen data for LLMs within a specific domain. For example, we can make use of those contents (e.g., journal articles) published subsequent to the release of the LLM.
To verify the effectiveness of our method, we conduct extensive experiments on various datasets, including WikiMIA, BookMIA (Shi et al., 2024), ArXivTection and BookTection (Duarte et al., 2024). The results demonstrate that our method can significantly improve the performance of existing methods based on scoring functions. For example, our method improves the AUC score of the best baseline method Min-k%, increasing it from 0.62 to 0.91 on WikiMIA under the OPT-6.7B model. Moreover, our method can also improve the TPR@5%FPR score of baseline methods. For example, our method improves the TPR@5%FPR score of the detection method using perplexity, increasing it from 0.10 to 0.81 on ArXivTection under the LLaMA-7B model.
Figure 1: Overview of Fine-tuned Score Deviation. To determine whether a text $X$ is in the pre-training data, our method first fine-tunes the pre-trained model with a few non-members and then measures the deviation distance of scores from the pre-trained model and fine-tuned model as a membership inference metric. If the deviation value is smaller than the threshold value, the text is likely in the pretraining data.
@@
  • We empirically show that FSD can improve the performance of existing detection methods based on scoring functions for detecting pretraining data, through extensive experiments conducted on various benchmark datasets with diverse large language models.
@@ 2 Background
Let $f$ be an autoregressive large language model (LLM) with trainable parameters $\bm{\theta}$ (e.g., LLaMA (Touvron et al., 2023a)) and $\mathcal{D}$ denotes the associated pretraining data, sampled from an underlying distribution $\mathcal{P}$. As model developers rarely provide detailed information about the pretraining dataset, we generally desire to identify if the LLM is trained on the given text for scientific and ethical concerns. Formally, the task objective is to learn a detector $h$ that can infer the membership of an arbitrary data point $\bm{x}$ in the dataset $\mathcal{D}$: $h(\bm{x},f_{\bm{\theta}})\rightarrow\{0,1\}$.
@@
3 Method: Fine-tuned Score Deviation
@@ 3.1 Motivation
Figures 2(a) and 2(b) present the deviation of perplexity distributions for members and non-members, throughout the fine-tuning on the non-member dataset $\widetilde{\mathcal{D}}$. The results show that unseen data in the pretraining tends to obtain a lower perplexity from the fine-tuned model compared to the pre-trained model. Contrarily, we observe that the deviation of perplexity distribution for members is negligible after the pre-trained model fine-tuning. The analysis indicates that fine-tuning with a few unseen data from a specific domain can decrease the likelihood-based scores of the LLM for other unknown non-members in the domain. The contrast in the score deviation resulting from fine-tuning allows for the distinction between members and non-members.
3.2 Method
@@
By way of the FSD score, we can obtain a clear distinction between members and non-members, establishing excellent performance in detecting pretraining data. To provide a straightforward view, we show in Figure 3 the score distribution between members and non-members using various scoring functions on WikiMIA (Shi et al., 2024). The results of ArXivTection (Duarte et al., 2024) are also presented in Appendix C.1. Our experiments validate that, compared to the perplexity and Min-kscores, our FSD score significantly increases the gap between non-members and members, and as a result, enables more effective pretraining data detection.
@@ 4.1 Experimental Setup
To verify the effectiveness of detection methods, we employ four common benchmark datasets for evaluations, including WikiMIA (Shi et al., 2024), ArXivTection (Duarte et al., 2024), BookTection (Duarte et al., 2024) and BookMIA (Shi et al., 2024). Previous works have demonstrated that model developers commonly use text content among those datasets for pre-training (Shi et al., 2024; Duarte et al., 2024; Ye et al., 2024). The datasets are provided by Hugging Face [^3] https://huggingface.co/datasets, and detailed information of datasets is presented in Appendix A.
@@
We use four detection methods based on scoring functions as our baselines for evaluating the performance of methods on various datasets under diverse models. Those methods employ specific metrics based on the likelihood, followed by a comparison with a preset threshold to identify the given text’s membership. Specifically, baseline methods include the example perplexity (Perplexity(Li, 2023), the ratio of example perplexity and zlib compression entropy (Zlib(Carlini et al., 2021), the ratio of the perplexity on the example before and after lowercasing (Lowercase(Carlini et al., 2021) and detecting pretraining example through outlier words with low probability (Min-k%(Shi et al., 2024).
@@
We evaluate the performance of detection methods for detecting pretraining data by measuring the following metrics: (1) AUC, the area under the receiver operating characteristic curve; (2) the true positive rate (TPR) when the false positive rate (FPR) of examples is 5% (TPR@5%FPR).
@@
Our approach involves constructing the non-member dataset and fine-tuning the base model. For constructing the non-member dataset, we randomly sample 30% of the data from the entire dataset and select all non-members from this subset as the constructed fine-tuning dataset. The remaining 70% of the dataset is used for testing. We employ LoRA (Hu et al., 2022) to fine-tune the base model with 3 epochs and a batch size of 8. We set the initial learning rate as 0.001 and drop it by cosine scheduling strategy. We conduct all experiments on NVIDIA L40 GPU and implement all methods with default parameters using PyTorch (Paszke et al., 2019).
@@
4.2 Main Results
Can FSD improve the performance of current scoring functions?
We compare the performance of detection methods on WikiMIA and ArXivTection datasets across various large language models. The detailed information of dataset split is shown in Appendix B.1. Our results in Table 1 show that the FSD significantly improves the performance of all baseline methods on both datasets across diverse models. For example, our method improves the AUC score compared to the best baseline method Min-k%, increasing it from 0.62 to 0.91 on WikiMIA dataset from the OPT-6.7B model. Similarly, it improves the AUC score from 0.76 to 0.86 on ArXivTection dataset from the LLaMA-7B model. Moreover, we show that our method also remarkably improves the TPR@5%FPR score of all baseline methods in Table 11 of Appendix C.2.
@@ 4.2 Main Results
To investigate the effect of varying the fine-tuning data size on the pretraining data detection, we compare the performance of the model fine-tuned with different-sized fine-tuned datasets. To construct fine-tuning datasets of varying sizes, we randomly sample varying amounts of non-members (0, 30, 50, 100, 150, 200, 250, 300) from the WikiMIA dataset as fine-tuning datasets. In addition, we sample 465 members and non-members from the dataset and construct a balanced test set of 930 examples.
Figure 4 presents the performance of FSD with various sizes of auxiliary datasets. The results show our method achieves better performance as the size of the fine-tuning dataset increases. Notably, our method is highly data-efficient, achieving dramatic improvements with only a small amount of non-members for fine-tuning. For example, FSD improves the AUC score of the perplexity-based method from 0.63 to 0.91, by leveraging only 100 non-member data for fine-tuning – a significant direct improvement of 44%. In summary, a few non-members are sufficient for FSD to improve the detection, demonstrating its practicality.
@@
We also verify the performance of baselines and our methods from different-sized LLaMA models (7B, 13B, 30B) on WikiMIA. In Table 2, our results demonstrate that our method is effective with different-sized models, and achieves remarkable performance from different-sized models. Notably, the AUC score of Lowercase slightly rises as the parameters of the LLaMA model increase. Moreover, additional results of the TPR@5%FPR score show a similar trend, which can be found in Appendix C.2.
(a) AUC
(b) TPR@5%FPR
@@
Recent works (Shi et al., 2024; Duarte et al., 2024) study the problem of copyrighted book detection in training data. Following previous works, we verify the effectiveness of detection methods for detecting excerpts from copyrighted books on BookMIA (Shi et al., 2024) and BookTection (Duarte et al., 2024). Specifically, we randomly sample 500 members and non-members from datasets, constructing a balanced validation set of 1,000 examples. The detailed information of datasets split is presented in Appendix B.2.
In Table 3, we compare the accuracy of our method and baselines for detecting suspicious books in pretraining data from the LLaMA-7B model. A salient observation is that our method significantly improves the accuracy of baseline methods for copyrighted book detection. For example, compared with baselines, our method reaches an accuracy of 98.6% on BookMIA using detection method Zlib, which marks a significant 71.8% improvement. We also present the AUC score with our method and baselines in Table 3, which consistently improves the detection capabilities of baseline methods. Our extensive experiments demonstrate the effectiveness of our method for copyrighted book detection.
@@ 5 DISCUSSION
The key step of our method is to fine-tune the pre-trained model using a few non-members. One may also ask: can a similar effect be achieved by utilizing members as the fine-tuning dataset? In this ablation, we separately sample members and non-members from WikiMIA to construct fine-tuning datasets(Mem, Non). In addition, we randomly sample data from WikiMIA as another fine-tuning dataset (All). The details of implementation are presented in Appendix B.3. To investigate the impact of different fine-tuning datasets on pretraining data detection, we fine-tune the LLaMA-7B model with each of the fine-tuning datasets individually.
Our results in Table 5 show that our method can improve the performance of baseline methods using members as the fine-tuning dataset. However, our method achieves inferior performance when using members for fine-tuning compared with non-members. Moreover, it is not realistic to construct a member dataset without accessing pretraining data in real-world scenarios. In addition, this is feasible for constructing non-members as a fine-tuning dataset based on the model release date and data creation timestamp. Overall, our method demonstrates superior performance when using non-members for fine-tuning, while ensuring applicability in real-world settings.
@@
Existing works generally construct benchmark datasets based on the LLM release date and data creation timestamp (Ye et al., 2024; Shi et al., 2024). For example, the WikiMIA dataset considers events occurring post-2023 as non-members. Recent works indicate evaluation results are suspect on benchmark datasets because they possibly sample members and non-members from different distributions (Duan et al., 2024; Das et al., 2024; Maini et al., 2024). We find the temporal shift between members and non-members in the WikiMIA dataset, which is shown in Appendix B.4. The issue shows that we can distinguish members and non-members with timestamps in the dataset. To eliminate the impact of temporal differences between members and non-members on evaluation, we implement two strategies to mitigate the temporal shift in the dataset: (1) removing timestamps in the dataset (Deletion), and (2) replacing the year of timestamps with 2023 in the dataset(Replacement). We conduct experiments with baselines and our method on the original WikiMIA dataset, Deletion and Replacement, respectively.
Our results in Table 5 show that the performance of baselines and our method are compromised when mitigating the temporal shift between members and non-members. This is likely because deleting or replacing a word will change the probability of the subsequent word, thereby perturbing the likelihood-based metric. In this setting, surprisingly, the results show that our method can also improve the performance of baseline methods. Though baseline methods yield results that are comparable to random guessing on the Replacement dataset, our method can also improve the AUC scores of the perplexity-based detection method, increasing it from 0.54 to 0.71. Overall, our method is effective even if there is no distribution difference between members and non-members. The TPR@5%FPR score of the experiment is presented in Appendix C.2.
@@
To expose LLMs to unseen data, we employ LoRA to fine-tune the pre-trained model. The results demonstrate that our method achieves impressive performance for pretraining data detection when fine-tuning with LoRA. However, can a similar effect be achieved using different fine-tuning methods? To this end, we apply AdaLoRA (Zhang et al., 2023), IA3 (Liu et al., 2022), and LoRA to fine-tune LLaMA-7B with WikiMIA, respectively. The details of the dataset in our experiment can be found in Appendix B.1.
@@
In Table 6, we report the AUC and TPR@5%FPR scores for pretraining data detection with our method and baseline methods. The results show that our method can improve the performance of baseline methods when using different fine-tuning methods. Although our FSD achieves inferior performance with AdaLoRA compared with IA3 and LoRA, it still significantly improves the performance of baseline methods. Overall, our method can be implemented with different fine-tuning methods and does not require a specific fine-tuning technique.
6 RELATED WORK

### v2

@@ Abstract
In the era of large language models (LLMs), detecting pretraining data has been increasingly important due to concerns about fair evaluation and ethical risks. Current methods differentiate members and non-members by designing scoring functions, like Perplexity and Min-k%. However, the diversity and complexity of training data magnifies the difficulty of distinguishing, leading to suboptimal performance in detecting pretraining data. In this paper, we first explore the benefits of unseen data, which can be easily collected after the release of the LLM. We find that the perplexities of LLMs shift differently for members and non-members, after fine-tuning with a small amount of previously unseen data. In light of this, we introduce a novel and effective method termed Fine-tuned Score Deviation (FSD), which improves the performance of current scoring functions for pretraining data detection. In particular, we propose to measure the deviation distance of current scores after fine-tuning on a small amount of unseen data within the same domain. In effect, using a few unseen data can largely decrease the scores of all non-members, leading to a larger deviation distance than members. Extensive experiments demonstrate the effectiveness of our method, significantly improving the AUC score on common benchmark datasets across various models. Our code is available at https://github.com/ml-stat-Sustech/Fine-tuned-Score-Deviation.
@@ 1 Introduction
In the literature, current works of detecting pretraining data primarily focus on designing scoring functions to differentiate members (i.e., seen data during pretraining) and non-members (unseen). For example, previous work shows that sequences from the training data tend to have lower perplexity (i.e., higher likelihood) than non-members (Li, 2023). Min-kleverages the kof tokens with minimum token probabilities of a text for detection, assuming that trained data tends to contain fewer outlier tokens (Shi et al., 2024). However, non-member data can obtain low perplexities by including frequent or repetitive texts, while members may contain rare tokens that result in high perplexities. This casts significant doubt on utilizing those scoring functions for detecting pretraining data. Consequently, this issue prompts us to present a preliminary attempt to enlarge the difference between members and non-members for pretraining data detection.
@@
Therefore, our key idea behind FSD is to enlarge the score deviation between members and non-members by exposing the LLM to a few non-members. This can be accomplished by measuring the deviation distance of current scores (See Figure 1), owing to the self-supervised fine-tuning on a few non-members. In effect, the fine-tuning largely decreases the scores of non-member data, resulting in more distinguishable seen and unseen data. In practical applications, it is easy to collect a small amount of unseen data for LLMs within a specific domain. For example, we can make use of those contents (e.g., journal articles) published subsequent to the release of the LLM.
To validate the effectiveness of our method, we conduct extensive experiments on various datasets, including WikiMIA, BookMIA (Shi et al., 2024), ArXivTection, BookTection (Duarte et al., 2024) and Pile (Maini et al., 2024). The results demonstrate that our method can significantly improve the performance of existing methods based on scoring functions. For example, our method improves the AUC score of the best baseline method Min-k%, increasing it from 0.62 to 0.91 on WikiMIA under the OPT-6.7B model. Moreover, our method can also improve the TPR@5%FPR score of baseline methods. For example, our method improves the TPR@5%FPR score of the detection method using perplexity, increasing it from 0.10 to 0.81 on ArXivTection under the LLaMA-7B model.
Figure 1: Overview of Fine-tuned Score Deviation. Our method first fine-tunes the pre-trained model with a few non-members and then measures the deviation distance of scores from the pre-trained and fine-tuned models as a membership inference metric. If the deviation value is smaller than the threshold value, the text $X$ is likely in the pretraining data.
@@
  • We empirically show that FSD can improve the performance of existing methods based on scoring functions for pretraining data detection, through extensive experiments conducted on various benchmark datasets with diverse LLMs.
@@ 2 Background
Pretraining data detection is an instance of membership inference attacks (MIAs) (Shokri et al., 2017), and can help identify data contamination in the pretraining corpus (Shi et al., 2024). Let $f$ be an autoregressive large language model (LLM) with trainable parameters $\bm{\theta}$ (e.g., LLaMA (Touvron et al., 2023a)) and $\mathcal{D}$ denotes the associated pretraining data, sampled from an underlying distribution $\mathcal{P}$. As model developers rarely provide detailed information about the pretraining dataset, we generally desire to identify if the LLM is trained on the given text for scientific and ethical concerns. Formally, the task objective is to learn a detector $h$ that can infer the membership of an arbitrary data point $\bm{x}$ in the dataset $\mathcal{D}$: $h(\bm{x},f_{\bm{\theta}})\rightarrow\{0,1\}$.
@@
3 Method
@@ 3.1 Motivation
Figures 2(a) and 2(b) present the deviation of perplexity distributions for members and non-members, throughout the fine-tuning on the non-member dataset $\widetilde{\mathcal{D}}$. The results show that unseen data in the pretraining tend to obtain a lower perplexity from the fine-tuned model than the pre-trained model. In contrast, we observe that the deviation of the perplexity distribution for members is negligible after fine-tuning. The analysis indicates that fine-tuning with a few unseen data from a specific domain can decrease the likelihood-based scores of the LLM for other unknown non-members in the domain. The contrast in the score deviation resulting from fine-tuning allows for the distinction between members and non-members.
3.2 FSD: Fine-tuned Score Deviation
@@
By way of the FSD score, we can obtain a clear distinction between members and non-members, establishing excellent performance in detecting pretraining data. To provide a straightforward view, we show in Figure 3 the score distribution between members and non-members using various scoring functions on WikiMIA (Shi et al., 2024). The results of ArXivTection (Duarte et al., 2024) are also presented in Appendix D.1. Our experiments validate that, compared to the perplexity and Min-kscores, our FSD score significantly increases the gap between non-members and members, and as a result, enables more effective pretraining data detection.
@@ 4.1 Experimental Setup
To verify the effectiveness of detection methods, we employ four common benchmark datasets for evaluations, including WikiMIA (Shi et al., 2024), ArXivTection (Duarte et al., 2024), BookTection (Duarte et al., 2024) BookMIA (Shi et al., 2024) and Pile (Maini et al., 2024). Previous works have demonstrated that model developers commonly use text content among those datasets for pre-training (Shi et al., 2024; Duarte et al., 2024; Ye et al., 2024). The datasets are provided by Hugging Face [^3] https://huggingface.co/datasets, and detailed information of datasets is presented in Appendix B.
@@
We use four detection methods based on scoring functions as baselines to evaluate the performance of different methods across various datasets and models. Those methods employ specific metrics based on the likelihood, followed by a comparison with a preset threshold to identify the membership of the given text. Specifically, baseline methods include the example perplexity (Perplexity(Li, 2023), the ratio of example perplexity and zlib compression entropy (Zlib(Carlini et al., 2021), the ratio of the perplexity on the example before and after lowercasing (Lowercase(Carlini et al., 2021) and detecting pretraining example through outlier words with low probability (Min-k%(Shi et al., 2024).
@@
We evaluate the performance of different detection methods for detecting pre-training data from large language models, by measuring the following metrics: (1) AUC, the area under the receiver operating characteristic curve; (2) the true positive rate (TPR) when the false positive rate (FPR) of the examples is 5% (TPR@5%FPR).
@@
Our approach involves constructing the non-member dataset to fine-tune the base model. For constructing the non-member dataset, we randomly sample 30% of the data from the entire dataset and select all non-members from this subset as the constructed fine-tuning dataset. The remaining 70% of the dataset is used for testing. We employ LoRA (Hu et al., 2022) to fine-tune the base model with 3 epochs and a batch size of 8. We set the initial learning rate as 0.001 and drop it by cosine scheduling strategy. We conduct all experiments on NVIDIA L40 GPU and implement all methods with default parameters using PyTorch (Paszke et al., 2019).
4.2 Main Results
Can FSD improve the performance of current scoring functions?
We compare the performance of detection methods on WikiMIA and ArXivTection datasets across various LLMs. The details of the dataset split are shown in Appendix C.1. The results in Table 1 show that the FSD significantly improves the performance of all baselines on both datasets across diverse models. For example, our method improves the AUC score compared to the best baseline method Min-k%, increasing it from 0.62 to 0.91 on the WikiMIA dataset under the OPT-6.7B model. Similarly, it improves the AUC score from 0.76 to 0.86 on the ArXivTection dataset under the LLaMA-7B model. Moreover, our method improves the TPR@5%FPR score of all baselines, which can be found in Appendix D.2. We also present the results on different subsets of the Pile dataset under the Pythia-6.9B model. Table 2 shows the average AUC scores of baselines and our method over the 20 subsets of Pile, demonstrating that our method achieves superior performance on the Pile dataset. In addition, the detailed results of all subsets of the Pile dataset are provided in Appendix D.3.
@@ 4.2 Main Results
Table 2: The average AUC score of baselines and our method from the Pythia-6.9B over 20 subsets of the Pile dataset. Base and +Ours respectively denote the baseline methods and our method.
Method Perplexity Lowercase Zlib Min-k%
Base +Ours Base +Ours Base +Ours Base +Ours
Pile 0.503 0.625 0.519 0.566 0.507 0.624 0.515 0.600
@@
To investigate the effect of varying the fine-tuning data size on the pretraining data detection, we compare the performance of our method using different-sized fine-tuned datasets for fine-tuning LLaMA-7B. To construct fine-tuning datasets of varying sizes, we randomly sample varying amounts of non-members (0, 30, 50, 100, 150, 200, 250, 300) from the WikiMIA dataset as fine-tuning datasets. In addition, we construct a balanced test set of 930 examples for evaluation.
Figure 4 presents the performance of FSD with various sizes of auxiliary datasets. The results show our method achieves better performance as the size of the fine-tuning dataset increases. Notably, our method is highly data-efficient, achieving dramatic improvements with only a small amount of non-members for fine-tuning. For example, FSD improves the AUC score of the perplexity-based method from 0.63 to 0.91, by leveraging only 100 non-member data for fine-tuning – a significant direct improvement of 44%. In summary, a few non-members are sufficient for FSD to improve the detection, demonstrating its practicality. In addition, we also evaluate our method on the BookC2 subset of the Pile dataset under the Pythia-6.9B model. The results show a similar trend, which can be found in Appendix D.3.
@@
We also verify the performance of baselines and our methods from different-sized LLaMA models (7B, 13B, 30B) on WikiMIA. In Table 3, our results demonstrate that our method is effective with different-sized models, and achieves remarkable performance from different-sized models. Notably, the AUC score of Lowercase slightly rises as the parameters of the LLaMA model increase. Moreover, additional results of the TPR@5%FPR score show a similar trend, which can be found in Appendix D.2.
@@
Recent works (Shi et al., 2024; Duarte et al., 2024) study the problem of copyrighted book detection in training data. Following previous works, we verify the effectiveness of detection methods for detecting excerpts from copyrighted books on BookMIA (Shi et al., 2024) and BookTection (Duarte et al., 2024). Specifically, we randomly sample 500 members and non-members from datasets, constructing a balanced validation set of 1,000 examples. The detailed information of datasets split is presented in Appendix C.2.
In Table 4, we compare the accuracy of our method and baselines for detecting suspicious books in pretraining data from the LLaMA-7B model. A salient observation is that our method significantly improves the accuracy of baseline methods for copyrighted book detection. For example, compared with baselines, our method reaches an accuracy of 98.6% on BookMIA using detection method Zlib, which marks a significant 71.8% improvement. We also present the AUC score with our method and baselines, which consistently improves the detection capabilities of baseline methods. Our extensive experiments demonstrate the effectiveness of our method for copyrighted book detection.
@@ 5 DISCUSSION
The key step of our method is to fine-tune the pre-trained model using a few non-members. One may also ask: can a similar effect be achieved by utilizing members as the fine-tuning dataset? In this ablation, we separately sample members and non-members from WikiMIA to construct fine-tuning datasets(Mem, Non). In addition, we randomly sample data from WikiMIA as another fine-tuning dataset (All). The details of implementation are presented in Appendix C.3. To investigate the impact of different fine-tuning datasets on pretraining data detection, we fine-tune the LLaMA-7B model with each fine-tuning dataset.
Our results in Table 6 show that our method can improve the performance of baseline methods using members as the fine-tuning dataset. However, our method performs inferior when using members for fine-tuning compared with non-members. Moreover, it is not realistic to construct a member dataset without accessing pretraining data in real-world scenarios. In addition, this is feasible for constructing non-members as a fine-tuning dataset based on the model release date and data creation timestamp. Overall, our method achieves superior performance when using non-members for fine-tuning, while ensuring applicability in real-world settings. We also investigate the performance of our method when fine-tuning using data from an unrelated domain. The results in Appendix D.3 show that our method can improve the performance when fine-tuning data from a mix of domains.
@@
Existing works generally construct benchmark datasets based on the LLM release date and data creation timestamp (Ye et al., 2024; Shi et al., 2024). For example, the WikiMIA dataset considers events occurring post-2023 as non-members. Recent works indicate evaluation results are suspect on benchmark datasets because they possibly sample members and non-members from different distributions (Duan et al., 2024; Das et al., 2024; Maini et al., 2024). We find the temporal shift between members and non-members in the WikiMIA dataset, which is shown in Appendix C.4. The issue shows that we may distinguish members and non-members with timestamps in the dataset. To eliminate the impact of temporal differences between members and non-members on evaluation, we implement two strategies to mitigate the temporal shift in the dataset: (1) removing timestamps in the dataset (Deletion), and (2) replacing the year of timestamps with 2023 in the dataset (Replacement). We conduct experiments with baselines and our method on the original WikiMIA dataset, Deletion and Replacement, respectively.
Our results in Table 6 show that our method also improves the performance of baseline methods when mitigating the temporal shift between members and non-members. In this setting, the performance of baselines and our method is reduced, as deleting or replacing a word will change the probability of the subsequent word, thereby perturbing the likelihood-based metric. Though baseline methods yield results comparable to random guessing on the Replacement dataset, our method improves the AUC scores of the perplexity-based method, increasing it from 0.54 to 0.71 on Replacement. Overall, our method is effective even if there is no distribution difference between members and non-members. The TPR@5%FPR score of the experiment is presented in Appendix D.2.
@@
To expose LLMs to unseen data, we employ LoRA to fine-tune the pre-trained model. The results demonstrate that our method achieves impressive performance for pretraining data detection when fine-tuning with LoRA. However, can a similar effect be achieved using different fine-tuning methods? To this end, we investigate the impact of fine-tuning methods on performance by applying AdaLoRA (Zhang et al., 2023), IA3 (Liu et al., 2022), and LoRA to fine-tune LLaMA-7B with WikiMIA, respectively.
@@
In Table 7, we report the AUC and TPR@5%FPR scores for pretraining data detection for our method and baselines. The results show that our method improves the performance of baselines when using different fine-tuning methods. Although our FSD achieves inferior performance with AdaLoRA compared with IA3 and LoRA, it still improves the performance of baseline methods. Our method can be implemented with different fine-tuning methods and does not require a specific fine-tuning technique. In addition, we also conduct experiments to explore the impact of different fine-tuning parameters on the performance of our method. The results show that our method is insensitive to LoRA rank and the number of fine-tuning epochs, which are presented in Appendix D.3.
6 CONCLUSION
Limitations
Our method requires to collect a few examples that belong to the same domain but are not involved in the training. Generally, we can utilize the data content published after the release of the LLM. Therefore, our method is applicable for detecting benchmarks or copyrighted resources in a specific domain (e.g., math tests, magazines). The diversity of the test set may make it challenging to construct an effective auxiliary dataset of unseen data. In addition, our method requires fine-tuning on a few non-member data, so the effectiveness of the proposed score might be affected by the data quality of non-members.
Ethical Statement
Our work focuses on pretraining data detection from large language models. The proposed methodology aims to address issues involving data contamination or copyright infringement. In addition, our method can be used to identify potential privacy leakage risks and ensure the safety of LLMs, aligning with established ethical standards for content moderation. Regarding data access, the evaluated datasets we employed in our work come from prior research and do not involve personal privacy information.
Acknowledgement
This research is supported by the Shenzhen Fundamental Research Program (Grant No. JCYJ20230807091809020) and the SUSTech-NUS Joint Research Program. Bingyi Jing is supported in part by the National Natural Science Foundation of China under grant 12371290. We gratefully acknowledge the support of the Center for Computational Science and Engineering at the Southern University of Science and Technology.
Appendix A RELATED WORK